2. A string vibrates with a certain fundamental (n 1) frequency. Which action will increase the frequency of the fundamental the most: a) cutting the length by one- half, b) reducing its mass by two-thirds, or c) doubling the tension? 3. In the experiment, how is the tension in the string determined?
Answer #1

2.

\nu=\frac {v}{2L}=\frac {1}{2L}\sqrt {\frac {T}{\mu}}

Cutting the length by one-half would increase the frequency of the fundamental the most.